I am desperately trying to help my niece with homework while her parents are out of town. I am not a math guru. Here is the word problem:
Joe is 2 years older than his brother. The sum of their ages is 12. How old is his brother?
She knows the answers are ages 5 and 7 (5+7=12) but I do not know how to help her figure out the mathematical equation.
Can you PLEASE HELP????

Answer by ankor@dixie-net.com(22740)

Let j = Joe's age
Let b = his brother's age
;
Write an equation for each statement
:
"Joe is 2 years older than his brother."
j = b + 2
:
"The sum of their ages is 12."
j + b = 12
In first equation we know that j = (b+2), replace j with (b+2) in the above equation
(b+2) + b = 12
add the b's
2b + 2 = 12
subtract 2 from both sides
2b = 12 - 2
2b = 10
Divide both sides by 2
b= 10/2
b = 5 yrs is brothers age.
:
Use the 1st equation to find Joes age
:
Check your solutions in the 2nd equatiion